Aaron is 12 years old. He is 3 times as old as his brother. How old will Aaron be when he is twice as old?
Answer by richwmiller(17219)

Twice as old as whom? himself? then he will be 24.
his brother?
his brother is 4
12=3*x
4=x
2+y=2*(4+y)
12+y=8+2y
4=y
Aaron will be twice as old as his brother in four years
Aaron will be 16 and his brother will be 8